Among other applications, oxytocin is used to induce labor. An IV dose of oxytocin can start at 0.002 units/min and then be increased by 0.001 units every 15 min until contractions become well established. How many hours will it take for a dose of 0.02 units/min to be reached? (not 5 hours)
Answer #1

Starting unit 0.002 units/min

Increasing 0.001 units every 15 mins

To reach 0.02 the 0.002 should be increased by 0.018 ( 0.02-0.002)

So total times required the increase = 0.018/0.001 = 18

So total time taken = 18* 15 minutes = 270 mts

= 270/60 = 4.5 hours